OpenStudy (anonymous):

Meno and Anytus both worked olive oil presses near Athens, Greece in 410 BCE. Anytus was able to produce three liters more than twice the oil that Meno produced in a morning. In fact, Anytus' output was only five liters less than three times Meno's output. The equation below models the relation between their rates of production, where x is the number of liters of oil that Meno could press. 2x + 3 = 3x − 5 How many liters of oil could Anytus press each morning?

hartnn (hartnn):

can u find 'x' from 2x + 3 = 3x − 5 which is Meno's output ?

hartnn (hartnn):

then Anyus output will be just 2x+3

OpenStudy (anonymous):

menos is 3x-5

hartnn (hartnn):

nopes, according to equation setup, Meno's output is taken as 'x' litres.

hartnn (hartnn):

both sides = 2x + 3 = 3x − 5 = anytus output 'x' is Menos output

hartnn (hartnn):

so, first find 'x'

OpenStudy (anonymous):

i dont know how to ok people i teird of every question find x i dont know how too

hartnn (hartnn):

2x + 3 = 3x − 5 add 5 to both sides, what u get ?

OpenStudy (anonymous):

umm i already got this wrong and i really dont feel like go over but thnxs anyway

OpenStudy (anonymous):

On a navigational chart of the Gulf coast of Florida, Panama City has coordinates (3, 20). A line running from Panama City to Tampa has a slope of −\[\frac{ 5 }{ 4 }\] What is the equation of the line? A. 5x − 4y = −65 B. 5x + 4y = 95 C. 4x + 5y = 112 D. 4x − 5y = −88

OpenStudy (anonymous):

candyland, if you don't let people teach you, you won't learn.

OpenStudy (anonymous):

true fine but i got like 50 more problems tonight

hartnn (hartnn):

m=5/4, (x1,y1) = (3,20) Use the equation, y-y1 = m (x-x1) plug in the values, what u get ?

OpenStudy (anonymous):

And if you let us teach you, you can do them really quickly yourself, instead of asking openstudy for all the questions You're rushing so much that you're getting all your questions wrong because you resorted to guessing.

OpenStudy (anonymous):

y-20=5/4(x-3)

OpenStudy (anonymous):

Learn, and you can do it quickly and correctly, okay?

OpenStudy (anonymous):

is that right hartnn

hartnn (hartnn):

yes, now you have to rearrange, to bring only constants on right, can u ?

OpenStudy (anonymous):

y-20=5/4x-3.75

hartnn (hartnn):

1)Add 20 to both sides 2)subtract 5/4 x from both sides 3)multiply by 4 on both sides and you get the answer.

OpenStudy (anonymous):

...wait, subtract 5/4 x from both sides?

hartnn (hartnn):

to bring x term on left side.

OpenStudy (anonymous):

yea i dont get that part i thought u divide

OpenStudy (anonymous):

Oh I see. I didn't look at the choices, sorry.

OpenStudy (anonymous):

umm im a bit confused what do i subract 5/4 from

OpenStudy (anonymous):

If you were trying to isolate x, then yes, it would be divide. But here, you're just moving the 5/4x to the other side.

hartnn (hartnn):

y-20=5/4x-3.75 y=5/4x-3.75+20 subtract 5/4 x from both sides, y-5/4 x = -3.75+20 got that ^ ?

OpenStudy (anonymous):

Both sides. So that you cancel out the right side.

OpenStudy (anonymous):

ohhhh ok so then add 5/4

hartnn (hartnn):

why ? then multiply 4 on both sides...

OpenStudy (anonymous):

where did u get four from its a fraction of 5/4

hartnn (hartnn):

yes, to convert fraction 5/4 into integer, i multiply both sides by 4...

OpenStudy (anonymous):

yesw i guess

OpenStudy (anonymous):

....y'know, we could've just used the general equation of a line.

OpenStudy (anonymous):

(psst although, if you got this far, don't forget to multiply both sides by -1 to get the actual answer....)